OpenStudy (anonymous):

What are the exact solutions of x2 − 3x − 7 = 0?

hartnn (hartnn):

u know quadratic formula ?

OpenStudy (anonymous):

uh mayybbbee \[\frac{ -b+-\sqrt{b ^{2}}-4ac }{ 2a }\]

hartnn (hartnn):

yes:) can u find a,b,c here ?

hartnn (hartnn):

and to be precise : the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

a=1 b=-3 c=-7?

hartnn (hartnn):

yes, now just put them in formula.

OpenStudy (anonymous):

ok hold on :)

OpenStudy (anonymous):

ok it would be \[x=\frac{ 3+-\sqrt{37} }{ 2}\]

OpenStudy (anonymous):

?

hartnn (hartnn):

yes, but when u solve that exact values are 4.5414 and -1.5414

OpenStudy (anonymous):

okay :) thank you! again! haha